# Parallax: A Compiler for Neutral Atom Quantum Computers under Hardware Constraints

### v1

@@ I Introduction
Quantum computing holds the promise to speed up computational tasks in diverse sectors, including scientific computing, high-performance computing, and machine learning [33, 34]. In the past, superconducting qubits have been at the forefront of quantum technology due to early breakthroughs that leverage our knowledge in classical silicon-based technologies [4]. Nevertheless, in recent years, neutral atom quantum computers have emerged as a promising alternative [39, 15].
Neutral atom systems have capabilities such as executing multi-qubit gates directly, allowing for application-specific qubit topologies, and the ability to move atoms to reconfigure qubit positions dynamically [47, 29, 18, 6, 9]. Additionally, the homogeneous nature of qubits and their ability for long-range interactions present a pathway to more scalable quantum computing solutions [46, 23, 44].
@@ I-A Brief Background of Quantum Technology
Atoms, once cooled in a magneto-optical trap (MOT), are trapped by an acousto-optic deflector (AOD) and a spatial light modulator (SLM) via the polarizing beam splitter (PBS) (Fig. 2[22, 23, 18, 37]. The SLM creates an array of static optical tweezer traps in a specified configuration, while the AOD provides a mobile array of optical traps, a feature central to the design principles of Parallax. After loading, the SLM displays a grid of atoms, each at a user-defined location. Operations on the qubit-atoms are conducted using specific lasers, and post-computation, a fluorescence-sensing camera reads out the atom array [1, 18].
@@
Gate Implementations. Computation is performed via different sets of lasers interacting with the atoms in the AOD/SLM arrays. Raman transitions, which involve pulses from lasers to transition an atom between the |$\ket{0}$ and |$\ket{1}$ hyperfine states, are used to perform the U3 gate with arbitrary rotation [18]. One of the main benefits of neutral atom systems is the ability to connect distant atoms using their Rydberg states. When an atom is highly excited, its outermost electron occupies a Rydberg state, meaning it has an enlarged orbit that allows interactions with nearby atoms. Local Rydberg lasers can be used to excite individual atoms [20]. These long-range atom interactions are used to implement CZ gates on atoms within the interaction radius of each other [23, 44]. While this allows for interactions between atoms, it also results in the Rydberg blockade effect: If one atom is in a Rydberg state, neighboring atoms within a certain radius cannot be similarly excited if they are not involved in a gate with that atom. This blockade effect restricts parallelism since any atoms in the blockade radius cannot themselves perform multi-qubit gates (though they can still perform single-atom Raman transitions). This blockade radius is typically 2.5×2.5$\times$2.5 × the interaction radius [42]. See Fig.3b(a) for an example of how Rydberg atoms interact with blockades.
@@
First, each atom has a minimum separation distance constraint, whereby it must never be closer to any other atom by some hardware-dependent distance [47, 41]. Second, the AOD grid is comprised of rows and columns of traps. AOD rows/columns cannot cross over each other, meaning the relative order of the rows and columns must be maintained for the duration of the program [9]. This is due to the fact that the AOD traps might interfere with one another if they are moved across each other. Lastly, qubits on an AOD row/column have to move in tandem [41, 9]. This means if one row has two atoms on it, and one atom is moved up by $x$ units, the other atom is moved up by $x$ as well.
@@ III Experimental Methodology
Experimental Setup. Parallax’s evaluation is conducted via a simulator that emulates the hardware characteristics of real neutral atom testbed systems. Note that while a physical 256-qubit neutral atom system does exist (QuEra Aquila [47]), it only supports analog computation and does not yet support a programmable AOD. Thus, to evaluate Parallax, we developed a simulator using real hardware parameters (see Table II). T1 and T2 times represent the decoherence speed for hyperfine atomic states; we get the decoherence error rates by inputting these times and each benchmark’s runtime into exponential decay functions. Note: if atoms are lost during a simulated shot because of atom collision or trap escape (modeled as being part of $T1$ error in our simulation), these atoms are replenished between physical shots as needed. Thus, the impact of atom loss is only on the error rates of the circuits. We use two different hardware simulations for our experimentation. We use QuEra’s 256-qubit Aquila system [47], which has a 16$\times$16 atom grid, for our main results, and Atom’s upcoming 1,225-qubit system [36], which has a 35$\times$35 atom grid, to evaluate scaling and parallelization. We use 20 AOD rows and columns as the default configuration. We also ablate it in the next section and show that 20 provides the best results.

### v2

@@ I Introduction
Quantum computing holds the promise to speed up computational tasks in diverse sectors, including scientific computing, high-performance computing, and machine learning [42, 43]. In the past, superconducting qubits have been at the forefront of quantum technology due to early breakthroughs that leverage our knowledge in classical silicon-based technologies [4]. Nevertheless, in recent years, neutral atom quantum computers have emerged as a promising alternative [51, 19, 50].
Neutral atom systems have capabilities such as executing multi-qubit gates directly, allowing for application-specific qubit topologies, and the ability to move atoms to reconfigure qubit positions dynamically [41, 59, 36, 23, 7, 11]. Additionally, the homogeneous nature of qubits and their ability for long-range interactions present a pathway to more scalable quantum computing solutions [58, 30, 56, isenhower2010demonstration].
@@ I-A Brief Background of Quantum Technology
Atoms, once cooled in a magneto-optical trap (MOT), are trapped by an acousto-optic deflector (AOD) and a spatial light modulator (SLM) via the polarizing beam splitter (PBS) (Fig. 2[29, 30, 23, 46, 6, 15]. The SLM creates an array of static optical tweezer traps [41, 49, 21, 14] in a specified configuration, while the AOD provides a mobile array of optical traps, a feature central to the design principles of Parallax [48, 27]. After loading, the SLM displays a grid of atoms, each at a user-defined location. Operations on the qubit-atoms are conducted using specific lasers, and post-computation, a fluorescence-sensing camera reads out the atom array [1, 50].
@@
Gate Implementations. Computation is performed via different sets of lasers interacting with the atoms in the AOD/SLM arrays. Raman transitions, which involve pulses from lasers to transition an atom between the |$\ket{0}$ and |$\ket{1}$ hyperfine states, are used to perform the U3 gate with arbitrary rotation [23]. One of the main benefits of neutral atom systems is the ability to connect distant atoms using their Rydberg states. When an atom is highly excited, its outermost electron occupies a Rydberg state, meaning it has an enlarged orbit that allows interactions with nearby atoms. Local Rydberg lasers can be used to excite individual atoms [26]. These long-range atom interactions are used to implement CZ gates on atoms within the interaction radius of each other [30, 56, isenhower2010demonstration]. While this allows for interactions between atoms, it also results in the Rydberg blockade effect: If one atom is in a Rydberg state, neighboring atoms within a certain radius cannot be similarly excited if they are not involved in a gate with that atom. This blockade effect restricts parallelism since any atoms in the blockade radius cannot themselves perform multi-qubit gates (though they can still perform single-atom Raman transitions). This blockade radius is typically 2.5×2.5$\times$2.5 × the interaction radius [54]. See Fig.3b(a) for an example of how Rydberg atoms interact with blockades.
@@
First, each atom has a minimum separation distance constraint, whereby it must never be closer to any other atom by some hardware-dependent distance [59, 24, 9]. Second, the AOD grid is comprised of rows and columns of traps. AOD rows/columns cannot cross over each other, meaning the relative order of the rows and columns must be maintained for the duration of the program [11]. This is due to the fact that the AOD traps might interfere with one another if they are moved across each other. Lastly, qubits on an AOD row/column have to move in tandem [53, 11]. This means if one row has two atoms on it, and one atom is moved up by $x$ units, the other atom is moved up by $x$ as well.
@@ III Experimental Methodology
Experimental Setup. Parallax’s evaluation is conducted via a simulator that emulates the hardware characteristics of real neutral atom testbed systems. Note that while a physical 256-qubit neutral atom system does exist (QuEra Aquila [59]), it only supports analog computation and does not yet support a programmable AOD. Thus, to evaluate Parallax, we developed a simulator using real hardware parameters (see Table II). T1 and T2 times represent the decoherence speed for hyperfine atomic states; we get the decoherence error rates by inputting these times and each benchmark’s runtime into exponential decay functions. Note: if atoms are lost during a simulated shot because of atom collision or trap escape (modeled as being part of $T1$ error in our simulation), these atoms are replenished between physical shots as needed. Thus, the impact of atom loss is only on the error rates of the circuits. We use two different hardware simulations for our experimentation. We use QuEra’s 256-qubit Aquila system [59], which has a 16$\times$16 atom grid, for our main results, and Atom’s 1,225-qubit system [37, 45], which has a 35$\times$35 atom grid, to evaluate scaling and parallelization. We use 20 AOD rows and columns as the default configuration. We also ablate it in the next section and show that 20 provides the best results.